# PoU: Proof-of-Use to Counter Tool-Call Hacking in DeepResearch Agents

### v1

@@ 4.6.3. Tool-Call Behavior Analysis
To assess how expanding the tool set changes agent behavior, we trained Deepresearcher+ with one more tool (Knowledge Graph) from scratch under the same data and hyperparameters as the original Deepresearcher, then uniformly sampled 1,0001{,}000 queries from seven test sets and counted tool-call usage across three categories: Web Search (Web), Knowledge Graph (KG), and Web Browser (Browser). Table 4 summarizes the call ratios.
Table 4. Tool-call ratio (%) over 1,0001{,}000 sampled queries from seven test sets. Columns are Web Search (Web), Knowledge Graph(KG), and Web Browser (Browser). Deepresearcher has no KG tool (0.0).
Model Web KG Browser
Deepresearcher 88.4 - 11.6
Deepresearcher+ 94.6 2.2 3.2
PoU 51.8 27.5 20.7
The results in Table 4 reveal distinct behavioral patterns across models. When trained with only the format and answer rewards (as in the original Deepresearcher setup), the agent tends to overuse available tools once the tool set expands. After adding the Knowledge Graph interface, Deepresearcher+ exhibits a pronounced bias toward redundant Web Search calls (94.6%) and a near disappearance of Web Browser usage (3.2%). This behavior indicates a collapse of tool specialization: although the browser is logically designed to serve as a follow-up step for Web Search—retrieving detailed content from previously retrieved URLs—the model largely ignores this expected hierarchy and instead repeatedly queries the Web Search proxy. Such uncontrolled escalation of retrieval actions suggests severe Tool-call Hacking when training lacks explicit evidence-usage constraints.
In contrast, the PoU-trained agent demonstrates a markedly more balanced tool distribution. Each proxy—Web Search, Knowledge Graph, and Browser—maintains a complementary role (51.8%, 27.5%, 20.7%, respectively), implying that the model actively schedules heterogeneous retrieval utilities based on their strengths: the Web Search proxy for broad discovery, the Knowledge Graph for exploratory search, and the Browser for fine-grained verification. This balanced behavior highlights that the PoU reward design effectively regularizes tool usage, encouraging agents to retrieve, reason, and cite evidence in a coordinated manner rather than relying on repetitive search actions.
4.6.4. Extreme Ablation
To further examine robustness and causal grounding, we replaced every <tool_response> with the token “content”, removing all semantics while keeping structure.
Model Behavior under ”content” responses
PoU <think>”No evidence; Try other tools.” $\rightarrow$ <helpful>no</helpful> $\rightarrow$ <ref>null</ref> $\rightarrow$ <tool_call> KGSearch(q)(q).
$\Rightarrow$ Detects missing evidence and switches tools/parameters.
[…]
$\Rightarrow$ Hallucinates evidence and continues fictitious reasoning.
Table 5. Extreme ablation test where all tool responses are replaced by “content”. Only PoU detects missing evidence and adapts its strategy.
@@ 5. Conclusion
In this work we identified tool‑call hacking as a pervasive failure mode in multi‑source retrieval‑augmented agents and introduced Proof‑of‑Use (PoU)—a contract‑driven, evidence‑grounded RL framework that enforces step‑wise alignment between retrieved evidence, intermediate reasoning, and final answers. In PoU this contract contains explicit helpfulness/citation declarations, a citation‑format reward, reference‑linked perturbation sensitivity, and an answer–citation alignment objective, coupled with selective unmasking of actually cited snippets during training. Across diverse in‑domain and out‑of‑domain QA benchmarks, PoU yields consistent gains in both F1 and LLM‑as‑judge scores over strong deep‑research baselines, while exhibiting balanced and purposeful tool routing and robustness under domain even tools shifts. Limitations include reliance on an external LLM judge for alignment scoring, and a current focus on retrieval tools rather than procedural tools. By the ablation studies, we clearly demonstrate the rationality and necessity of each reward design. Future work will explore learned or weakly supervised judges, adaptive perturbation curricula, cost‑ and latency‑aware routing, extensions to calculators/coders/simulators, and robustness against adversarial or privacy‑sensitive tool outputs. Overall, PoU offers a simple and auditable path toward trustworthy multi‑source reasoning by making evidence use not only observable but optimizable.
Ethical Use of Data and Informed Consent
This study does not involve human participants or personally identifiable information. All datasets used (e.g., HotpotQA, PubMed, and other publicly available corpora) are released under open research licenses and comply with their respective terms of use. All web content accessed complied with robots.txt and site terms, and was used solely for non-commercial research. No additional ethical approval was required.

### v2

@@ 4.5.3. Tool-Call Behavior Analysis
Building on Figure 1, which illustrates divergent tool-entropy dynamics between PoU and the baseline during training, we further examine how agents adapt their tool-call patterns under different task settings and tool environments. We uniformly sample 1,000 rollouts from each of seven open-domain test sets (HotpotQA, 2Wiki, NQ, TQ, …) and two out-of-tool-distribution (OOT) benchmarks (BioASQ, PQArefEval), and report the resulting tool-call ratio distribution across tools in Figure 4.
Figure 4. Tool-call ratio (%) distributions across models and settings. For visual clarity, values below 6% are not shown. For brevity, DeepR. denotes DeepResearcher.
With only the format and answer rewards (as in the original DeepResearcher setup), the agent tends to overuse available tools once the tool set expands. After trained with one more tool proxy, DeepResearcher+ exhibits a pronounced bias toward redundant Web Search calls (94.6%) and a near disappearance of Web Browser usage (3.2%). This behavior indicates a collapse of tool specialization: despite the intended search–browse hierarchy, the agent fails to differentiate tool roles and repeatedly queries the Web Search proxy. This indicates not merely a preference imbalance, but a failure to learn functional differentiation among tools. Similarly, under OOT and NT settings, DeepResearcher maintains a highly polarized and stable tool-call distribution, reflecting a rigid tool usage pattern that persists despite changes in task relevance. From a training perspective, this behavior arises because outcome-level rewards dominate optimization. Once a particular tool consistently yields non-zero rewards during early training, the policy is reinforced to repeatedly invoke it, while alternative tools that lack step-level corrective signals are gradually suppressed. As a result, the agent converges to a brittle retrieval strategy that fails to adapt under tool-set or distribution shifts. Such retrieval actions suggests severe Tool-call Hacking of normal format-answer supervision again.
In contrast, the PoU agent exhibits a more balanced tool-call distribution. Importantly, beyond the robustness observed in OOT and NT settings, this balance reflects a second emergent property: a non-static and context-sensitive tool allocation that varies across tasks and tool environments, indicating flexible adaptation of heterogeneous retrieval utilities.
4.5.4. Extreme Case
To expose tool-call hacking and evaluate resistance to spurious tool response, we intervene in tool-call by randomly replacing the content of the corresponding <tool_response> with the token ‘‘content’’, while preserving the response structure. Due to space constraints, Table 4 only presents the <think> component of a single reasoning step under spurious tool responses. Nevertheless, this simplified illustration faithfully reflects the behavior patterns we consistently observe in full rollouts. In practice, when the tool response content is corrupted while preserving its structure, DeepResearcher often produces seemingly coherent intermediate reasoning that implicitly assumes the existence of valid evidence, despite the absence of any usable information in the tool output. In contrast, PoU explicitly recognizes the lack of supporting evidence at the step level, marks the tool call as unhelpful, and proactively resorts to alternative tools. This contrast is robust across queries and steps, and the extreme case shown here serves as a minimal yet representative example of the underlying phenomenon.
Table 4. An Extreme Case
Observation:<tool_response> content </tool_response>
Method
DeepResearcher <think>
I found the information about “xxx”.
Now I will check “xxx”.
</think>
PoU <think>
<helpful>no</helpful>
<ref>null</ref>
No relevant information found.
I will try another tool.
</think>
@@ 5. Conclusion
In this work, we identify tool-call hacking as a subtle yet critical failure mode in RL-trained deep research agents, arising from the weak observability of evidence usage under outcome-dominated supervision. To address this issue, we propose Proof-of-Use (PoU), an evidence-grounded training framework that explicitly enforces causal dependencies between retrieval, reasoning, and final answers. Extensive experiments across diverse tasks and tool environments demonstrate that PoU effectively mitigates tool-call hacking and yields strong performance. Beyond this, PoU exhibits notable emergent properties, including robustness and adaptive tool-usage behaviors under domain and tool shifts, despite not being explicitly optimized for tool adaptation.